OpenStudy (anonymous):

What is the factorization of the expression below? 4x2 - 81

OpenStudy (btaylor):

using the difference of squares: \[(a^2-b^2) = (a+b)(a-b)\] Think of \[4x^2 = 2x \times 2x\] and \[81 = 9 \times 9\] If 4x^2=a^2 and 81=b^2, you get a=2x and b=9 So, factored would be (2x+9)(2x-9)